OpenStudy (anonymous):

100 meters divided by 10.49 seconds?

OpenStudy (anonymous):

rounding 100 / 10.5 = 9.5 meters per second?

OpenStudy (anonymous):

\[\frac{100m}{10.49s} = \frac{100}{10.49} \frac{m}{s}\]

OpenStudy (anonymous):

well. that would be 1ft because 1 meter = 3.28ft and 3.28/3.28 = 1

OpenStudy (anonymous):

5,280ft equals what

OpenStudy (anonymous):

1 mile